# Redshift leverage for the search of GRB neutrinos affected by quantum properties of spacetime

### v1

@@ V Avenues for refining the strategy of analysis
In this Section we discuss and investigate some observations which could be used to refine the strategy of analysis. We believe that some of the observations reported in this Section deserve being considered in setting up in-vacuo-dispersion analyses of future neutrino data, even though the quantitative aspects of this Section should be assessed with caution since they were produced in a second phase of our investigations after seeing the data, and might therefore be affected by the sources of bias that are well known for unblind analyses. In particular, some of the observations reported in this Section are at least in part inspired by the fact that we saw that among our four GRB-neutrino candidates there are three that are described particularly well according to Eq. (1), and by our perception that the procedure used in Sec. IV to estimate the significance of our findings did not “benefit” sufficiently from the noteworthy properties of those three best GRB-neutrino candidates.
@@ V.2 Improved correlation
One challenge whose handling might require further methodological refinements in future studies concerns the choice of an appropriate statistic quantifying the agreement between the properties of the found GRB-neutrino candidates and Eq. (1). Investigations of this aspect should take into account the fact that the time window for in-vacuo dispersion searches of GRB neutrinos will always be rather large, even if, as one might hope, the $\eta$ range on which one would focus could become narrower with the accrual of more data. A relatively small uncertainty on $\eta$ still translates (especially at high neutrino energies and high redshift) into a rather large time window, large enough for a tangible probability of background contamination of the GRB-neutrino candidates. These concerns are even more severe if one allows for non-systematic fuzzy in-vacuo dispersion effects of the type mentioned here briefly in Sec. I (and discussed in more detail in, e.g., Refs. [16, 17, 24]), since in that case searches of GRB-neutrino candidates would always have to explore a range of $\eta$ with width at least $\delta\eta$.
@@ V.4 Scanning a range of values of $\eta$
For simplicity we illustrate and discuss the strategy of analysis based on this alternative statistic applying it directly to the data already analyzed in this paper. For every given value of $\eta$ in our interval of interest $[12.7,30.7]$, we find the number $N(\eta)$ of GRB-neutrino candidates which are directionally compatible in the usual sense ($\mathcal{S}\geq 1/4\pi$) and satisfy Eq. (1) for that specific value of $\eta$ within two standard uncertainties in the neutrino energy. We then use simulated data to estimate the probability that the null hypothesis would accidentally produce at least $N(\eta)$ such GRB-neutrino candidates, finding a corresponding $p$-value $p(\eta)$. The resulting “$p$-curve” is reproduced here as Fig. 3.
@@
The statistical significance of our $p$-curve can be characterized estimating how frequently the minimum of the analogous $p$-curves built out of simulated data happens to be less than or equal to the minimum value $p(20.8)=0.004$ attained in Fig. 3. The resulting $p$-value is $P_{ns}=0.012$.
@@ VI Outlook
The essential equivalence of the $p$-values $P_{[26]}$, characterizing the statistical significance of the previous analysis [26], and $P_{r}$, found here in Sec. IV and characterizing in the same way the significance of our current analysis, shows that the new, redshift-based selection criteria for GRB-neutrino candidates proposed in Sec. I are competitive with the ones adopted in Ref. [26]. This means that the improvement in data quality provided by the knowledge of the GRB redshifts is more than sufficient to compensate for the loss of most of the GRB dataset (we measure the redshift of about 1/10 of observed GRBs). We feel that this is our main result, as it effectively opens up a new avenue, complementary to the approach pursued in Ref. [26], to investigate in-vacuo dispersion with neutrinos.
@@
We expect that the redshft-based strategy articulated in the previous sections will prove to be the best starting point for any future refinements, but we do not exclude that a suitable way may still be found for including GRBs whose redshift has not been measured (of course, from the perspective of our approach, somehow estimating the redshift of those GRBs with a suitably large uncertainty). Such a refinement of our approach could prove even more valuable in a few years, as more IceCube data is accrued and KM3NeT starts observing high-energy neutrinos.
At least equally important for the analysis of future IceCube and KM3NeT data would be to remove the other key limitation of our selection criteria, the restriction to shower-event neutrinos. This in principle could be achieved by incorporating the known [42] large and non-Gaussian probability distribution (whose peak lies sizably above the deposited energy) that relates the primary neutrino energy to the deposited energy in track events.
@@ Acknowledgements.
We thank Anastasia Tsvetkova for guidance on the available GRB catalogues and resources, which was very valuable in the early stages of our work on Appendix A. G.A.-C. and G.G. are grateful for financial support by the Programme STAR Plus, funded by Federico II University and Compagnia di San Paolo, and by the MIUR, PRIN 2017 grant 20179ZF5KS. G.R.’s work on this project was supported by the National Science Centre grant 2019/33/B/ST2/00050. G.D.’s work on this project was supported by the Research Council of Norway, project number 301718. G.F.’s work on this project was supported by “Fondazione Angelo Della Riccia” and “The Foundation Blanceflor”. This work also falls within the scopes of the COST Action CA23130 “Bridging high and low energies in search of quantum gravity” and the COST Action CA18108 ”Quantum Gravity phenomenology in the Multi-Messenger era”.
@@ Appendix A Known-redshift GRB data
The following table contains the GRB data used in our analysis. We include all GRBs observed from January 2010 to December 2022 for which a definite redshift determination is available in the literature. We did not include GRBs with only upper and/or lower bounds on redshift. For the selection of the sample and the redshift values $z$ we relied mostly on Greiner’s catalogue [33], referred to in our list as MP. Searching the literature for other GRBs whose redshift was measured, we found 24 which are not acknowledged as such in the summary table of Ref. [33]. For these GRBs we provide references to the relevant works and GCNs in the column “$z$ ref.” of our table.
@@ Appendix B GRBs relevant for Figure 1
As already stressed in the main text, our analysis relied on the latest HESE neutrino data release by IceCube  [31]. We considered only shower neutrinos, for which the energy is approximately equal to the observed-deposited energy [40, 42] (whereas the energy of track neutrinos is poorly known).
@@
The neutrino with lowest energy among those relevant for Figure 1 has a deposited energy of 55.3 TeV and is associated to GRB110503A [76], a long GRB (T90=10.0 s) with measured redshift of 1.613 [77].
The violet point in Fig. 1 is the one that was already taken into account in the analysis of [26], with deposited energy of 64.0 TeV, and is associated to GRB111229A [78], a long GRB (T90=25.4 s) with measured redshift of 1.38 [79].
The neutrino with deposited energy of 302.2 TeV which is relevant for Fig. 1 is associated to GRB120923A [80], a long GRB (T90=27.2 s) with measured redshift of 7.8 [81].
The highest energy neutrino relevant for Fig. 1 (also known as “big bird” [82]) has deposited energy of 2075.0 TeV and is associated with GRB120909A [83], a long GRB (T90=112.1 s) with measured redshift of 3.93 [84].

### v2

@@ V Avenues for refining the strategy of analysis
In this Section we discuss and investigate some observations which could be used to refine the strategy of analysis. We believe that some of the observations reported in this Section deserve being considered in setting up in-vacuo-dispersion analyses of future neutrino data, even though some of the quantitative aspects of this Section should be assessed with caution since they were produced in a second phase of our investigations after seeing the data, and might therefore be affected by the sources of bias that are well known for unblind analyses. In particular, some of the observations reported in this Section are at least in part inspired by the fact that we saw that among our four GRB-neutrino candidates there are three that are described particularly well according to Eq. (1), and by our perception that the procedure used in Sec. IV to estimate the significance of our findings did not “benefit” sufficiently from the noteworthy properties of those three best GRB-neutrino candidates.
@@ V.2 Improved correlation
One challenge whose handling might require further methodological refinements in future studies concerns the choice of an appropriate statistic quantifying the agreement between the properties of the found GRB-neutrino candidates and Eq. (1). Investigations of this aspect should take into account the fact that the time window for in-vacuo dispersion searches of GRB neutrinos will always be rather large, even if, as one might hope, the $\eta$ range on which one would focus could become narrower with the accrual of more data. A relatively small uncertainty on $\eta$ still translates (especially at high neutrino energies and high redshift) into a rather large time window, large enough for a tangible probability of background contamination of the GRB-neutrino candidates. These concerns are even more severe if one allows for non-systematic, fuzzy in-vacuo dispersion effects of the type mentioned here briefly in Sec. I (and discussed in more detail in, e.g., Refs. [18, 19, 26]), since in that case searches of GRB-neutrino candidates would always have to explore a range of $\eta$ with width at least $\delta\eta$.
@@ V.4 Scanning a range of values of $\eta$
For simplicity we illustrate and discuss the strategy of analysis based on this alternative statistic applying it directly to the data already analyzed in this paper. For every given value of $\eta$ in our interval of interest $[12.7,30.7]$, we find the number $N(\eta)$ of GRB-neutrino candidates which are directionally compatible in the usual sense ($\mathcal{S}\geq 1/4\pi$) and satisfy Eq. (1) for that specific value of $\eta$ within two standard uncertainties in the neutrino energy. We then use simulated data to estimate the probability that the null hypothesis would accidentally produce at least $N(\eta)$ such GRB-neutrino candidates, finding a corresponding (local) $p$-value $p(\eta)$. The resulting “$p$-curve” is reproduced here as Fig. 3.
@@
The statistical significance of our $p$-curve can be characterized estimating how frequently the minimum of the analogous $p$-curves built out of simulated data happens to be less than or equal to the minimum value $p(20.8)=0.004$ attained in Fig. 3. The resulting (global) $p$-value is $P_{ns}=0.012$.
The fact that $P_{ns}$ and $P_{r}$, resulting from considerably different strategies of analysis, are of the same order of magnitude suggests that both values are in fact genuine characterizations of the statistical significance of our results, i.e., that they are not too dependent on the specific properties of the corresponding statistics.
@@ VI Outlook
The essential equivalence of the $p$-values $P_{[34]}$, characterizing the statistical significance of the previous analysis [34], and $P_{r}$, found here in Sec. IV and characterizing in the same way the significance of our current analysis, shows that the new, redshift-based selection criteria for GRB-neutrino candidates proposed in Sec. I are competitive with the ones adopted in Ref. [34]. This means that the improvement in data quality provided by the knowledge of the GRB redshifts is more than sufficient to compensate for the loss of most of the GRB dataset (we measure the redshift of about 1/10 of observed GRBs). We feel that this is our main result, as it effectively opens up a new avenue, complementary to the approach pursued in Ref. [34], to investigate in-vacuo dispersion for neutrinos.
@@
We expect that the redshift-based strategy articulated in the previous sections will prove to be the best starting point for any future refinements, but we do not exclude that a suitable way may still be found for including GRBs whose redshift has not been measured (of course, from the perspective of our approach, somehow estimating the redshift of those GRBs with a suitably large uncertainty). Such a refinement of our approach could prove even more valuable in a few years, as more IceCube and KM3NeT data is accrued.
At least equally important for the analysis of future IceCube and KM3NeT data would be to remove the other key limitation of our selection criteria, the restriction to shower-event neutrinos. This could in principle be achieved by incorporating into the analysis the known [42] large and non-Gaussian probability distribution (whose peak lies sizably above the visible energy) that relates the primary neutrino energy to the visible energy in track events.
@@ Acknowledgements.
We thank Anastasia Tsvetkova for guidance on the available GRB catalogues and resources, which was very valuable in the early stages of our work on Appendix A. G.A.-C. and G.G. are grateful for financial support by the Programme STAR Plus, funded by Federico II University and Compagnia di San Paolo, and by the MIUR, PRIN 2017 grant 20179ZF5KS. G.R.’s work on this project was supported by the National Science Centre grant 2019/33/B/ST2/00050. G.D.’s work on this project was supported by the Research Council of Norway, project number 301718, and by the Beatriu de Pinós programme (BP 2023). G.F.’s work on this project was supported by “Fondazione Angelo Della Riccia” and “The Foundation Blanceflor”. This work also falls within the scopes of the COST Action CA23130 “Bridging high and low energies in search of quantum gravity” and the COST Action CA18108 ”Quantum Gravity phenomenology in the Multi-Messenger era”.
@@ Appendix A Known-redshift GRB data
The following table contains the GRB data used in our analysis. We include all GRBs observed up to December 2022 for which a definite redshift determination is available in the literature. We did not include GRBs with only upper and/or lower bounds on redshift. For the selection of the sample and the redshift values $z$ we relied mostly on Greiner’s catalogue [45], referred to in our list as MP. Searching the literature for other GRBs whose redshift was measured, we found 45 which are not acknowledged as such in the summary table of [45]. For these GRBs we provide references to the relevant works and GCNs in the column “$z$ ref.” of our table.
@@ Appendix B GRBs relevant for Figure 1
As already stressed in the main text, our analysis relied on the latest HESE neutrino data release by IceCube [39]. We considered only shower neutrinos, for which the energy is approximately equal to the visible energy [41, 42] (whereas the energy of track neutrinos is poorly known).
@@
[…]